OpenStudy (anonymous):

Solve. 2x^2 – 3x + 6 = 0

OpenStudy (anonymous):

complex solutions for this one use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]with \[a=2,b=-3,c=6\]

OpenStudy (anonymous):

get \[x=\frac{3\pm\sqrt{-39}}{4}\]

OpenStudy (anonymous):

Solve. x^2 + 4x – 4 = 0

OpenStudy (anonymous):

I just tried solving this. But i got the wrong answer. Can you help me again?):

OpenStudy (anonymous):

real solutions for this one

OpenStudy (anonymous):

easiest to write \[x^2+4x=4\] \[(x+2)^2=4+2^2+8\] \[x+2=\pm\sqrt{8}=\pm2\sqrt{2}\] \[x=-2\pm2\sqrt{2}\]

OpenStudy (anonymous):

typo there. second line should be \[(x+2)^2=4+2^2=8\]

OpenStudy (anonymous):

yes a typo!